OpenStudy (anonymous):

Solve: log (2x + 1) - log x = log 4

OpenStudy (ash2326):

\[\ \log (2x + 1) - \log x = \log 4\] We know that \[\log a-\log b=\log \frac ab\] Applying this here, we get \[\log \frac{2x+1} {x}=\log 4\] Now we take anti log both sides \[\frac{2x+1} {x}= 4\] Can you solve for x now?

OpenStudy (anonymous):

Can you keep going just in case. o; I'm super paranoid. Lool.

OpenStudy (ash2326):

Give it a try, if you get stuck I'll help you

OpenStudy (anonymous):

I got 1/2

OpenStudy (ash2326):

Good, that's right:D

OpenStudy (anonymous):

Tysm C: